# Derandomizing Multivariate Polynomial Factoring for Low Degree Factors

### v1

@@ Abstract
    quasipolynomial time, for $\mathcal{C}$ being sparse polynomials,
@@
Result 2 and 3 were already shown by Kumar, Ramanathan, and Saptharishi with a different proof and their time complexities necessarily depend on black-box PITs for a related bigger class $\mathcal{C}^{\prime}$. Our complexities vary on whether the input is given as a blackbox or whitebox.
@@ 1 Introduction
The problem of multivariate polynomial factorization asks to find the unique factorization of a given polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ as a product of distinct irreducible polynomials over $\mathbb{F}$. The problem reduces to univariate polynomial factorization over the same field, for which a deterministic polynomial time algorithm is known over the field $\mathbb{Q}$. The complexity of multivariate factorization depends on the representation of input and output polynomials. If we use dense representation (where all the coefficients are listed including the zero coefficients), deterministic polynomial time algorithms for multivariate factoring are known [Kal85a]. If we use sparse representation (where only the nonzero coefficients are listed), only randomized polynomial time (in the total sparsity of input polynomial and the output factors) algorithms are known [vzGK85, KT90]. There are other standard representations like arithmetic circuits, and blackbox models (that gives the evaluations of the polynomial at any point, but the internal structure of the computation is hidden). Randomized polynomial time factorization algorithms are known in these models due to the classic results of Kaltofen [Kal89] and Kaltofen and Trager [KT90].
Towards derandomization of special cases of multivariate factoring, we ask the following.
Question 1 (Promise factoring[vzGK85]).
Let $f=\prod_{i=1}^{m}{g_{i}}^{e_{i}}$, where each $g_{i}$ is an irreducible polynomial with its sparsity at most $s$. Design a better-than-exponential time deterministic algorithm in $s,n,d$, to output $g_{i}$.
To our surprise, we do not know a deterministic subexponential-time algorithm even for the special case of Question 1, when the given blackbox computes the product of just two irreducible sparse polynomials.
@@
Question 2.
Design a better-than-exponential time deterministic algorithm that outputs all the sparse irreducible factors of a sparse polynomial.
Multivariate polynomial factoring has various applications, such as low-degree testing [AS03], constructions of pseudorandom generators for low-degree polynomials [Bog05, DGV24], computational algebraic geometry [HW00] and many more. Blackbox multivariate polynomial factorization is extensively used in arithmetic circuit reconstruction [Shp07, Sin16], and polynomial equivalence testing [Kay, Kay12, RR19]. Algebraic hardness vs randomness [KI03] results crucially use multivariate factorization.
@@
    Now often Hensel lifting is used to lift the projected factors back to the real factors. In our setting, we avoid Hensel lifting and argue that we can use interpolation instead. This interpolation trick was first used in the randomized black-box factoring algorithm of Kaltofen and Trager [KT90].
@@
Given an arithmetic circuit $C$, polynomial identity testing (PIT) asks to test if $C$ computes the zero polynomial. Randomization in factorization algorithms mostly stem from the fact that these algorithms use PIT as a subroutine. Further, Kopparty, Saraf and Shpilka [KSS15] showed that derandomization of white-box and blackbox multivariate circuit factoring reduces to derandomization of polynomial identity testing of arithmetic circuits in white-box and blackbox settings respectively. However, we do not know if sparse factorization reduces to sparse PIT or constant-depth arithmetic circuit PIT (the algorithms of [KSS15] reduce to general arithmetic circuit PIT). Special cases of depth-4 polynomial identity testing are related to questions about sparse polynomial factorization [Gup14, Vol17, BV22]. Recently, there has been some progress on these questions by [KRS24, KRSV24]. Earlier works of Volkovich [Vol15, Vol17] made progress on several special cases of sparse multivariate factoring. Shpilka and Volkovich [SV10] proved white-box/black-box factoring of multilinear polynomials in a class $\mathcal{C}$ is equivalent to white-box/black-box derandomization of PIT of class $\mathcal{C}$.
@@
In a factorization algorithm, we may want to check if a candidate factor is truly a factor via divisibility testing. It asks to test if a polynomial $g({\boldsymbol{z}})$ divides a polynomial $f({\boldsymbol{z}})$. Forbes [For15] showed that the divisibility testing question can be efficiently reduced to an instance of a PIT question of a model that relates to both $f$ and $g$, see Lemma 2.8. Currently, we do not know any deterministic polynomial time algorithm even when $g$ and $f$ are both sparse polynomials. When $f$ is a sparse polynomial and $g$ is a linear polynomial, the problem reduces to polynomial identity testing of any-order read-once oblivious branching programs (ROABPs), see Section 3.4. We do not know a deterministic polynomial time algorithm, even for testing if a quadratic polynomial $g$ divides a sparse polynomial.
@@
For testing irreducibility of multivariate polynomials, one can use some of the effective versions of Hilbert’s irreducibility theorem [vzGK85, Kal85b, Kal95]. These results can be also seen as an effective version of a classical theorem of Bertini in algebraic geometry. A multivariate irreducible polynomial $f(x,z_{1},\ldots,z_{n})$ may become reducible if we project the variables to make it univariate, but irreducibility is preserved if we project it in a way to make it a bivariate polynomial. Unfortunately, finding such a projection in deterministic polynomial time is hard. We do not know how to find such irreducibility preserving projections even for sparse polynomials, though we know efficient hitting set generators for them.
@@ 1.1 Our results
For a specific class $\mathcal{C}$, it suffices now to consider the complexity of the above two points that we already described before. For example for $\mathcal{C}$ being polynomials that have only constant degree factors, this yields a polynomial time algorithm for computing the constant degree factors, because PIT is in polynomial time and we can skip the divisibility test (Theorem 3.1). For $\mathcal{C}$ being sparse polynomials, we get a quasipolynomial-time algorithm (Corollary 3.3), and for $\mathcal{C}$ being polynomials computed by constant-depth circuits, we get subexponential-time (Corollary 3.4).
The last two results were already shown by Kumar, Ramanathan and Saptharishi [KRS24] with a different technique, using Hensel lifting. However, there is a subtle difference. In our case, we maintain the setting. That is, when the input is given whitebox, we use whitebox algorithms, and similarly for blackbox. On the other hand, the factoring algorithm by Kumar, Ramanathan and Saptharishi [KRS24] requires blackbox PIT algorithms, even when the input is given in whitebox. This is due to PIT for the resultant polynomial where the unknown factor is involved.
In the above two results, this difference does not matter because we have the same complexity bounds for whitebox and blackbox algorithms there. An example where this difference matters are commutative read-once oblivious arithmetic branching programs ($\rm ROABP$). There is a polynomial time whitebox PIT algorithm for ROABPs [RS05], while the best-known blackbox PIT algorithm for commutative ROABPs runs in quasipolynomial time [GKS17, GG20]. Moreover, it follows from work of Forbes [For15] that the divisibility test for $\rm ROABP$s by linear polynomials reduces to a PIT for $\rm ROABP$s. Hence, we have again polynomial, resp. quasipolynomial time for the divisibility test in the whitebox, resp. blackbox setting. It follows that the linear factors of polynomials computed by commutative $\rm ROABP$s can be computed in polynomial time in the whitebox setting, whereas it takes quasipolynomial time in the blackbox setting (Corollary 3.8).
@@
  • PIT for $\mathcal{C}$,
  • divisibility tests of $\mathcal{C}$ by constant-degree polynomials,
@@
For $\mathcal{C}$ being sparse polynomials, the difficulty lies in the third point: only exponential-time algorithms are known to derandomize Hilbert Irreducibility Theorem for sparse polynomials. The hardness stems from the fact that a sparse polynomial may have both sparse and non-sparse irreducible factors. Hence, preserving irreducibility for sparse polynomials will not preserve the factorization pattern, and therefore, it may be hard to get back the actual factor. Nevertheless, Theorem 4.2 pinpoints the challenge to compute the sparse factors of sparse polynomials (Corollary 4.4 and 4.5).
@@ 2 Preliminaries
We take $\mathbb{F}=\mathbb{Q}$ as the underlying field throughout the paper, although the results hold as well over fields with large characteristics.
@@
${\boldsymbol{z}}^{{\boldsymbol{e}}}=z_{1}^{e_{1}}\,z_{2}^{e_{2}}\,\cdots\,z_{n}^{e_{n}}\,.$
Its degree is $||{\boldsymbol{e}}||_{1}=\sum_{i=1}^{n}e_{i}$. The number of monomials of a polynomial $f$ with nonzero coefficient is called the sparsity of $f$ and is denoted by $\textup{sp}(f)$. The class of polynomials with sparsity $s$ is denoted by
@@
We define class $\partial\mathcal{C}$ as all the partial the derivatives of polynomials in $\mathcal{C}$,
@@
Let $x$ and ${\boldsymbol{z}}=(z_{1},\dots,z_{n})$ be variables and $f(x,{\boldsymbol{z}})$ be a $(n+1)$-variate polynomial. Then we can view $f$ as a univariate polynomial $f=\sum_{i}a_{i}({\boldsymbol{z}})\,x^{i}$ over $\mathbb{K}[x]$, where $\mathbb{K}=\mathbb{F}[{\boldsymbol{z}}]$. The $x$-degree of $f$ is denoted by $\deg_{x}(f)$. It is the highest degree of $x$ in $f$. Polynomial $f$ is called monic in $x$, if the coefficient $a_{d_{x}}({\boldsymbol{z}})$ is the constant $1$ polynomial, i.e. $a_{d_{x}}({\boldsymbol{z}})=1$, where $d_{x}=\deg_{x}(f)$.
An algorithm runs in subexponential time, if its running time on inputs of length $n$ and $\varepsilon>0$ is bounded by $2^{n^{\varepsilon}}$.
@@ 2.1 Computational problems and complexity measures
For the decision problems  ${\rm PIT}({\mathcal{P}})$ and ${\rm Div}({\mathcal{P}}/{\mathcal{Q}})$ there is an associated construction problem. In case of ${\rm PIT}({\mathcal{P}})$, a decision algorithm also yields an algorithm that computes a point ${\boldsymbol{a}}\in(\mathbb{F}\setminus\{0\})^{n}$ such that $p({\boldsymbol{a}})\not=0$, in case when $p\not\equiv 0$.
@@
In the blackbox case, let $H_{0}$ be the queries of the decision algorithm on the input of the zero-polynomial. Note that $H_{0}$ is a hitting set for the whole class ${\mathcal{P}}$. Hence, we can find ${\boldsymbol{a}}\in H_{0}$ in time $\textup{T}_{{\rm PIT}({\mathcal{P}})}$ such that $p({\boldsymbol{a}})\not=0$.
Still, some coordinates of ${\boldsymbol{a}}$ might be 00. In this case, we shift ${\boldsymbol{a}}$: Let $t$ be a new variable and consider ${\boldsymbol{a}}+t=(a_{1}+t,a_{2}+t,\dots,a_{n}+t)$. Then $p({\boldsymbol{a}}+t)$ is a nonzero polynomial in one variable of degree $d$. Let $M=|a_{i}|$, where $a_{i}$ is the minimum coordinate of ${\boldsymbol{a}}$. Then all coordiantes of ${\boldsymbol{a}}+M+1$ are positive and there is a $t\in\{M+1,M+2,\dots,M+d+1$}, such that $p({\boldsymbol{a}}+t)\not=0$. For the time complexity to find the right $t$, we have to add $(d+1)$ evaluations of $p$.
In the whitebox case, one can search for ${\boldsymbol{a}}$ by assigning values successively to the variables and do kind of a self-reduction. For each variable, one tries at most $d$ values from $\{1,2,\dots,d\}$ for a polynomial of degree $d$. If they all give 00, definitely $d+1$ works because it cannot be zero at $(d+1)$ many values. With $n$ variables, this amounts to $nd$ calls to the ${\rm PIT}$-decision algorithm. ∎
@@
  • $\mu(f)={\rm size}_{\rm ROABP}(f)$, the width of the smallest read read-once oblivious branching program (ROABP) that computes $f$.
@@
For example, when we skip the index $\mu$, we just refer to circuit size,

### v2

# arXiv:2411.17330v2[cs.CC] 27 Feb 2026

@@ Abstract
    quasi-polynomial time, for $\mathcal{C}$ being sparse polynomials,
@@
Result 2 and 3 were already shown by Kumar, Ramanathan, and Saptharishi with a different proof, and their time complexities necessarily depend on blackbox PITs for a related bigger class $\mathcal{C}^{\prime}$. Our complexities vary on whether the input is given as a blackbox or whitebox.
@@ 1 Introduction
The problem of multivariate polynomial factorization asks to find the unique factorization of a given polynomial $f\in\mathbb{F}[x_{1},\ldots,x_{n}]$ as a product of distinct irreducible polynomials over $\mathbb{F}$. The problem reduces to univariate polynomial factorization over the same field, for which a deterministic polynomial time algorithm is known over the field $\mathbb{Q}$. The complexity of multivariate factorization depends on the representation of input and output polynomials. If we use dense representation (where all the coefficients are listed including the zero coefficients), deterministic polynomial time algorithms for multivariate factoring are known [Kaltofen85]. If we use sparse representation (where only the nonzero coefficients are listed), only randomized polynomial time algorithms are known [von1985factoring, kaltofen1990computing]. There are other standard representations like arithmetic circuits, and blackbox models, where the evaluations of the polynomial at any point, but the internal structure of the computation is hidden. Randomized polynomial time factorization algorithms are known in these models due to the classic results of Kaltofen [kaltofen1989factorization] and Kaltofen and Trager [kaltofen1990computing]. Randomization is naturally required for these models, as the more basic question of polynomial identity testing (PIT) is not yet derandomized and multivariate circuit factorization algorithms use PIT. However, for various restricted arithmetic circuits we know deterministic (quasi-) polynomial- or subexponential-time PIT algorithms. Naturally, one can ask whether factorization can be derandomized for the corresponding circuit classes. In various problems such as special cases of arithmetic circuit reconstruction, we encounter multivariate factoring of structured or very restricted instances, where either the input polynomial or some or all factors of the input polynomial belong to a very restricted class of circuits. Towards derandomization of special cases of multivariate factoring, we focus on the following two goals.
Challenge 1 (Promise factoring [von1985factoring]).
Let $f=\prod_{i=1}^{m}{g_{i}}^{e_{i}}$, where each $g_{i}$ is an irreducible polynomial with its sparsity at most $s$. Design a deterministic polynomial-time algorithm to output $g_{i}$.
To our surprise, we do not know a deterministic polynomial-time algorithm even for the special case of Challenge 1, when the given blackbox computes the product of just two irreducible sparse polynomials.
@@
Challenge 2.
Given a sparse polynomial $f$, design a deterministic polynomial-time algorithm that outputs all the sparse irreducible factors of $f$.
Multivariate polynomial factoring has various applications, such as low-degree testing [arora2003improved], constructions of pseudorandom generators for low-degree polynomials [bogdanov2005pseudorandom, dwivedi2024optimal], algorithmic questions in computational algebraic geometry [huang2000extended] and many more. Blackbox multivariate polynomial factorization algorithms are extensively used in arithmetic circuit reconstruction [shpilka2007interpolation, sinha2016reconstruction], and polynomial equivalence testing [Kayal11, kayal2012affine, ramya2019linear]. There is a series of papers starting with [kabanets2003derandomizing] about algebraic hardness vs. randomness results that crucially rely on the closure of various algebraic complexity classes under factorization. For a general overview of multivariate factoring, we refer to the recent survey of Bhargav, Dwivedi, and Saxena [bhargav2025primer].
@@
    Now often Hensel lifting is used to lift the projected factors back to the real factors. In our setting, we avoid Hensel lifting and argue that we can use interpolation instead. This interpolation trick was first used in the randomized blackbox factoring algorithm of Kaltofen and Trager [kaltofen1990computing].
@@
Given an arithmetic circuit $C$, polynomial identity testing (PIT) asks to test if $C$ computes the zero polynomial. Randomization in factorization algorithms mostly stem from the fact that these algorithms use PIT as a subroutine. Further, Kopparty, Saraf and Shpilka [kopparty2015equivalence] showed that derandomization of whitebox and blackbox multivariate circuit factoring reduces to derandomization of polynomial identity testing of arithmetic circuits in whitebox and blackbox settings respectively. However, we do not know if sparse factorization reduces to sparse PIT (the algorithms of [kopparty2015equivalence] reduce to general arithmetic circuit PIT). Recently, there has been good progress on these questions by [kumar2024deterministic, kumar2024towards, bhattacharjee2025deterministic, bhattacharjee2025closure]. Earlier works of Volkovich [volkovich2015deterministically, volkovich2017some] and Bhargava, Saraf and Volkovich [bhargava2020deterministic] made progress on derandomization of several special cases of sparse multivariate factoring. Shpilka and Volkovich [shpilka2010relation] proved whitebox/blackbox factoring of multilinear polynomials in a class $\mathcal{C}$ is equivalent to whitebox/blackbox derandomization of PIT of class $\mathcal{C}$. For a comprehensive overview of derandomization questions related to multivariate polynomial factoring, we refer to the survey of Forbes and Shpilka [forbes2015complexity].
@@
In a factorization algorithm, we may want to check if a candidate factor is truly a factor via divisibility testing. It asks to test if a polynomial $g({\boldsymbol{z}})$ divides a polynomial $f({\boldsymbol{z}})$. Forbes [forbes2015deterministic] showed that the divisibility testing question can be efficiently reduced to an instance of a PIT question of a model that relates to both $f$ and $g$, see Lemma 2.8. Currently, we do not know any deterministic polynomial time algorithm even when $g$ and $f$ are both sparse polynomials. When $f$ is a sparse polynomial and $g$ is a linear polynomial, the problem reduces to polynomial identity testing of any-order read-once oblivious branching programs (ROABPs), see Section 3.4. We do not know a deterministic polynomial-time algorithm, even for testing if a quadratic polynomial $g$ divides a sparse polynomial.
@@
For testing irreducibility of multivariate polynomials, one can use some of the effective versions of Hilbert’s Irreducibility Theorem [von1985factoring, kaltofen1985effective, kaltofen1995effective]. These results can be also seen as an effective version of a classical theorem of Bertini in algebraic geometry. A multivariate irreducible polynomial $f(x,z_{1},z_{2},\ldots,z_{n})$ may become reducible if we project the variables to make it univariate, but irreducibility is preserved if we project it in a way to make it a bivariate polynomial. Unfortunately, finding such a projection in deterministic polynomial time is hard. We do not know how to find polynomial-sized irreducibility preserving projections even for sparse polynomials, though we know polynomial-sized hitting set generators for them. This remains an outstanding open question.
Special cases of multivariate factoring.
In many settings, one has a priori structural information about the factors of a given polynomial— for instance, that certain (or even all) factors are linear, have bounded degree, or more generally, are sparse. A classic source of such a structure is the theory of $U$-resultants due to Macaulay, where one encounters families of polynomials that decompose into linear forms. As another example, consider the standard arithmetization of a $3$-CNF formula: it naturally produces a polynomial presented as a product of clause polynomials, and each clause polynomial depends on at most three variables and is multilinear Restricted factorizations of this type have also been studied in the context of algebraization [impagliazzo2009axiomatic].
There are many prior works on special cases of multivariate polynomial factoring. Koiran and Ressyare [koiran2018orbits] gave three different randomized algorithms for testing, if a given polynomial $f$ as a blackbox can be completely factored into linear polynomials and output the factorization, if it exists. Medini and Shpilka [medini2021hitting] gave a deterministic polynomial-time algorithm in the blackbox setting for the case, where the polynomial can be factored into linearly independent linear polynomials.
Factorization of sparse multivariate polynomials is well-studied in symbolic computation. The classic work of Gathen and Kaltofen [von1985factoring] gives a randomized polynomial-time algorithm in the case when the input and output factors are sparse, and the number of irreducible factors is bounded. Volkovich [volkovich2015deterministically, volkovich2017some] made interesting progress on derandomization of some special cases of sparse multivariate factoring. Bhargava, Saraf, and Volkovich [bhargava2020deterministic] gave a deterministic polynomial time algorithm for factoring sparse polynomials with bounded individual degree. Special cases of depth-4 polynomial identity testing are related to questions about sparse polynomial factorization [gupta2014algebraic, volkovich2017some, bisht2022solving].
The sequence of papers [kumar2024deterministic, kumar2024towards, bhattacharjee2025deterministic] made significant progress on derandomization of factoring constant depth circuits. The recent breakthrough result by Bhattacharjee et al. [bhattacharjee2025closure] showed that all factors of constant depth circuits are computable by poly-sized constant depth circuits, and they gave a deterministic subexponential-time algorithm to compute those circuits.
@@ 1.1 Our results
  • PIT for the top-degree homogeneous component of polynomials in $\mathcal{C}$,
  • divisibility tests of the partial derivatives of $\mathcal{C}$ by constant-degree polynomials.
In many cases, class $\mathcal{C}$ is closed under taking homogeneous components and partial derivatives. Then the above two oracles simplify to
@@
For a specific class $\mathcal{C}$, it suffices now to consider the complexity of the above two points that we already described before. For example for $\mathcal{C}$ being polynomials that have only constant degree factors, this yields a polynomial time algorithm for computing the constant degree factors, because PIT is in polynomial time and we can skip the divisibility test (Theorem 3.1). For $\mathcal{C}$ being sparse polynomials, we get a quasi-polynomial-time algorithm (Corollary 3.3), and for $\mathcal{C}$ being polynomials computed by constant-depth circuits, we get subexponential-time (Corollary 3.4).
The last two results were already shown by Kumar, Ramanathan and Saptharishi [kumar2024deterministic] with a different technique, using Hensel lifting. However, there is a subtle difference. In our case, we maintain the setting. That is, when the input is given whitebox, we use whitebox algorithms, and similarly for blackbox. On the other hand, the factoring algorithm by Kumar, Ramanathan and Saptharishi [kumar2024deterministic] requires blackbox PIT algorithms, even when the input is given in whitebox. This is due to PIT for the resultant polynomial where the unknown factor is involved.
In the above two results, this difference does not matter because we have the same complexity bounds for whitebox and blackbox algorithms there. An example where this difference matters are commutative read-once oblivious arithmetic branching programs ($\rm ROABP$). There is a polynomial time whitebox PIT algorithm for ROABPs [raz2005deterministic], while the best-known blackbox PIT algorithm for commutative ROABPs runs in quasi-polynomial time [DBLP:journals/toc/GurjarKS17, guo2020improved]. Moreover, it follows from work of Forbes [forbes2015deterministic] that the divisibility test for $\rm ROABP$s by linear polynomials reduces to a PIT for $\rm ROABP$s. Hence, we have again polynomial, resp. quasi-polynomial time for the divisibility test in the whitebox, resp. blackbox setting. It follows that the linear factors of polynomials computed by commutative $\rm ROABP$s can be computed in polynomial time in the whitebox setting, whereas it takes quasi-polynomial time in the blackbox setting (Corollary 3.8).
@@
  • PIT for the top-degree homogeneous component of polynomials in $\mathcal{C}$,
  • divisibility tests of the partial derivatives of $\mathcal{C}$ by constant-degree polynomials,
@@
Again, if class $\mathcal{C}$ is closed under taking homogeneous components and partial derivatives, the first two items boil down to a PIT for $\mathcal{C}$ and divisibility tests of $\mathcal{C}$ by constant-degree polynomials.
For $\mathcal{C}$ being sparse polynomials, the difficulty lies in the third point: only subexponential-time algorithms are known to derandomize Hilbert Irreducibility Theorem for sparse polynomials [bhattacharjee2025closure]. The hardness stems from the fact that a sparse polynomial may have both sparse and non-sparse irreducible factors. Hence, preserving irreducibility for sparse polynomials will not preserve the factorization pattern, and therefore, it may be hard to get back the actual factor. Nevertheless, Theorem 4.2 pinpoints the challenge to compute the sparse factors of sparse polynomials (Corollary 4.4 and 4.5).
@@ 2 Preliminaries
We take $\mathbb{F}=\mathbb{Q}$ as the underlying field throughout the paper, although the results hold as well over fields with large characteristic.
@@
${\boldsymbol{z}}^{{\boldsymbol{e}}}=z_{1}^{e_{1}}\,z_{2}^{e_{2}}\,\cdots\,z_{n}^{e_{n}}\,.$ (2)
Its degree is ${\boldsymbol{e}}||_{1}=\sum_{i=1}^{n}e_{i}$. The number of monomials of a polynomial $f$ with nonzero coefficient is called the sparsity of $f$ and is denoted by $\textup{sp}(f)$. The class of polynomials with sparsity $s$ is denoted by
@@
We define class $\partial\mathcal{C}$ as all the partial derivatives of polynomials in $\mathcal{C}$,
@@
Let $x$ and ${\boldsymbol{z}}=(z_{1},z_{2},\dots,z_{n})$ be variables and $f(x,{\boldsymbol{z}})$ be a $(n+1)$-variate polynomial. Then we can view $f$ as a univariate polynomial $f=\sum_{i}a_{i}({\boldsymbol{z}})\,x^{i}$ over $\mathbb{K}[x]$, where $\mathbb{K}=\mathbb{F}[{\boldsymbol{z}}]$. The $x$-degree of $f$ is denoted by $\deg_{x}(f)$. It is the highest degree of $x$ in $f$. Polynomial $f$ is called monic in $x$, if the coefficient $a_{d_{x}}({\boldsymbol{z}})$ is the constant $1$ polynomial, i.e. $a_{d_{x}}({\boldsymbol{z}})=1$, where $d_{x}=\deg_{x}(f)$.
An algorithm runs in subexponential time, if its running time on inputs of length $n$ is bounded by $2^{O(n^{\varepsilon})}$, for any $\varepsilon>0$.
@@ 2.1 Computational problems and complexity measures
For the decision problems  ${\rm PIT}({\mathcal{P}})$ and ${\rm Div}({\mathcal{P}}/{\mathcal{Q}})$ there is an associated construction problem. In case of ${\rm PIT}({\mathcal{P}})$, a construction algorithm also yields a point ${\boldsymbol{a}}\in(\mathbb{F}\setminus\{0\})^{n}$ such that $p({\boldsymbol{a}})\not=0$, in case when $p\not\equiv 0$.
@@
In the blackbox case, let $H_{0}$ be the queries of the decision algorithm on input of the zero-polynomial. Note that $H_{0}$ is a hitting set for the whole class $\mathcal{P}$}. Hence, we can find ${\boldsymbol{a}}\in H_{0}$ in time $\textup{T}_{{\rm PIT}({\mathcal{P}})}$ such that $p({\boldsymbol{a}})\not=0$.
Still, some coordinates of $\boldsymbol{a}$} might be 0. In this case, we shift $\boldsymbol{a}$}: Let $t$ be a new variable and consider ${\boldsymbol{a}}+t=(a_{1}+t,a_{2}+t,\dots,a_{n}+t)$. Then $p({\boldsymbol{a}}+t)$ is a nonzero polynomial in one variable of degree $d$. Choose a set $S\subseteq\mathbb{F}-\{-a_{1},-a_{2},\dots,-a_{n}\}$ of size $|S|=d+1$. Then ${\boldsymbol{a}}+t$ has only non-zero coordinates, for all $t\in S$, and there is a $t\in S$ such that $p({\boldsymbol{a}}+t)\not=0$. For the time complexity to find the right $t$, we have to add $(d+1)$ evaluations of $p$.
In the whitebox case, one can search for $\boldsymbol{a}$} by assigning values successively to the variables and do kind of a self-reduction. For each variable, one tries at most $d$ values from $\{1,2,\dots,d\}$ for a polynomial of degree $d$. If they all give 0, definitely $d+1$ works because it cannot be zero at $(d+1)$ many values. With $n$ variables, this amounts to $nd$ calls to the $\rm PIT$}-decision algorithm. ∎
@@
  • $\mu(f)={\rm size}_{\rm ROABP}(f)$, the width of the smallest read-once oblivious branching program (ROABP) that computes $f$.
@@
When we skip the index $\mu$, we just refer to circuit size,